OpenStudy (anonymous):

@hartnn Write the sum using summation notation, assuming the suggested pattern continues. 5 - 15 + 45 - 135 + ...

hartnn (hartnn):

do you see any pattern in that?? like if you multiply/divide current term with a certain number, you get the next term??

OpenStudy (anonymous):

the cerrent term multiplied by 3

hartnn (hartnn):

if you multiply it by just a 3, then you will get all positive numbers, try multiplying by -3!

OpenStudy (anonymous):

so \[\sum_{n=0}^{\infty}5(-3)^n\]

OpenStudy (anonymous):

?

hartnn (hartnn):

so, using a summation notation in general, \(\Large \sum \limits_{n=0}^{n=\infty } f(n)\) oh you already got it!

OpenStudy (anonymous):

ok thanks!

hartnn (hartnn):

yes thats correct :)